Calculate the energy of (a) a photon of frequency 7.0×10^14 Hz , (b) a photon of wavelength 3.0×10^-.7 m.
1
Expert's answer
2020-11-10T07:08:13-0500

(a)

"E=h\nu=6.63 \sdot 10^{-34}\cdot 7\cdot 10^{14}=4.64\cdot 10^{-19}"J

(b)

"E=h\nu=h\frac{c}{\lambda}=6.63\cdot10^{-34}\cdot\frac{3\cdot 10^8}{3\cdot10^{-7}}=6.63\cdot10^{-19}" J